# Cascading Large Language Models for Salient Event Graph Generation

### v1

@@ 6 Conclusion
This study explored utilising LLMs to generate salient event relation graphs from news documents without relying on human annotations. We studied how the events generated by LLMs are compared to the traditional methods in terms of event saliency. We further demonstrated that CALLMSAE-generated graphs can serve as distant signals to fine-tune smaller models and outperform those based on CAEVO.
@@
Although we have tested many prompting methods and included several of the most effective ones in this paper, we have not explored all possible combinations due to the extensive volume of recent literature on prompt engineering. There might still exist combinations of prompts that could further improve performance. However, we are almost certain that any potential combinations, if they exist, are likely to be more complex and thus less efficient for building large-scale datasets. For example, we did not add demonstrations in graph generation because the code template is already quite lengthy. Adding more documents could potentially exceed the context windows of some LLMs, making it challenging for them to interpret the instructions effectively.
@@ Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2).
@@ Appendix B Saliency Features
| Write a summary of the document below using one paragraph. Document: """ Residents of British Columbia will receive a big cut in their income taxes on New Year’s Day, their second in six months. But far from winning applause, the province’s governing Liberal Party is experiencing a drop in public support. Tax cuts were a major promise in the campaign that brought the Liberals to office in the province in a landslide election win last May. The Liberals, led by Gordon Campbell, a former secondary school teacher, real-estate executive and three-term mayor of Vancouver, won all but two of the 79 seats in the western province’s legislative assembly. They defeated the left-leaning New Democrats, whose 10 years in office were marked by a growing public role in the economy and numerous tales of economic mismanagement. … """ Summary: | The Liberal Party government in British Columbia, led by Gordon Campbell, has implemented a significant cut in income taxes on New Year’s Day, fulfilling a campaign promise, but it has coincided with an economic downturn. Despite the tax cuts, the province is facing a large deficit due to falling tax revenues and increased demand for social services. To address this, the government plans to reduce the civil service by 28%, impose a three-year spending freeze on healthcare and education, and tighten benefits under the government-financed drug plan. However, many British Columbians are unhappy with these measures, and the Liberal Party’s approval rating has dropped significantly, amidst concerns about the return of political polarization in the province. |
@@
Appendix C Prompting Details
The total time cost for prompting Llama3 to construct the training data ( documents) is about hours (total wall-clock time of all the machines). We run the models on multiple machines with different specifications, including one with RTX , one with an A, and one with A.

### v2

@@ 6 Conclusion
This study explored utilising LLMs to generate salient event relation graphs from news documents without relying on human annotations. We studied how the events generated by LLMs are compared to the traditional methods in terms of event saliency. We further demonstrated that CALLMSAE-generated graphs can serve as distant signals to fine-tune smaller models and outperform those based on CAEVO. The CALLMSAE-generated event graphs, together with the human-annotated test set are collected as NYT-SEG.
@@
CALLMSAE is more demanding than CAEVO in terms of computational power and time. To generate the NYT-SEG, we spent wall clock hours in inference. More details about resource cost are disclosed in Appendix D.
Although we have tested many prompting methods and included several of the most effective ones in this paper, we have not explored all possible combinations due to the extensive volume of recent literature on prompt engineering. There might still exist combinations of prompts that could further improve performance Alzaid et al. (2024); Zhou et al. (2024). However, we are almost certain that any potential combinations, if they exist, are likely to be more complex and thus less efficient for building large-scale datasets. For example, we did not add demonstrations in graph generation because the code template is already quite lengthy. Adding more documents could potentially exceed the context windows of some LLMs, making it challenging for them to interpret the instructions effectively. The main goal of this work is to demonstrate the potential of LLM-based generation can help the data-demanding event graph generation task.
@@ Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2). Xingwei Tan was supported by the Warwick Chancellor’s International Scholarship. This work was conducted on the Sulis Tier-2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium.
@@ Appendix B Saliency Features
| Write a summary of the document below using one paragraph. Document: """ Residents of British Columbia will receive a big cut in their income taxes on New Year’s Day, their second in six months. But far from winning applause, the province’s governing Liberal Party is experiencing a drop in public support. Tax cuts were a major promise in the campaign that brought the Liberals to office in the province in a landslide election win last May. The Liberals, led by Gordon Campbell, a former secondary school teacher, real-estate executive and three-term mayor of Vancouver, won all but two of the 79 seats in the western province’s legislative assembly. They defeated the left-leaning New Democrats, whose 10 years in office were marked by a growing public role in the economy and numerous tales of economic mismanagement. With the Jan. 1 reductions, personal income tax rates for provincial taxes – which constitute a much larger share of overall taxes for Canadians than state taxes do for Americans – will have been chopped by an average of 25 percent, turning British Columbia from one of Canada’s most heavily taxed provinces to one with some of the lowest tax rates. … """ Summary: | The Liberal Party government in British Columbia, led by Gordon Campbell, has implemented a significant cut in income taxes on New Year’s Day, fulfilling a campaign promise, but it has coincided with an economic downturn. Despite the tax cuts, the province is facing a large deficit due to falling tax revenues and increased demand for social services. To address this, the government plans to reduce the civil service by 28%, impose a three-year spending freeze on healthcare and education, and tighten benefits under the government-financed drug plan. However, many British Columbians are unhappy with these measures, and the Liberal Party’s approval rating has dropped significantly, amidst concerns about the return of political polarization in the province. |
@@
Appendix C Comparison of Formats
In our preliminary experiments, we tested asking LLMs to generate event relations in JSON format (Table 10). When generating JSON response, we found LLMs often make format errors (e.g., using double quotation marks in the wrong place) which makes the responses difficult to parse, resulting in lower HGS. We also found the JSON format is not as flexible as the Python format which can integrate more information (e.g., explanations and definitions) as comments.
Appendix D Prompting Details
Due to limited resources, the sampling of the distant supervision dataset was conducted on multiple machines with different specifications, including one with RTX , one with an A, and one with A. The total time cost for prompting Llama3 to construct the training data ( documents) is about hours (the wall-clock time of all the machines combined).